OpenStudy (anonymous):

solve system equations by substitution x+y=4 -2x-2y=-8

OpenStudy (anonymous):

i came up with (3,1)

OpenStudy (anonymous):

is that correct

OpenStudy (anonymous):

in x+y = 4, minus the x from both sides, so y = 4-x, plug in 4-x into your bottom equation and solve, ... -2x -2(4-x) = -8

OpenStudy (anonymous):

once you have your x value, plug that into first equation to get your y

OpenStudy (anonymous):

is it (-4,8)

OpenStudy (anonymous):

no i still think its (1,3)

OpenStudy (anonymous):

the answer is the identity

OpenStudy (anonymous):

whats the identity

OpenStudy (anonymous):

it's when there is an infinite number of solutions, not just one when you solve the equations by substitution you end up with -8=-8 which is always true

OpenStudy (anonymous):

so is it (0,4) for my answer

OpenStudy (anonymous):

no the answer is any set of numbers that will work in the first equation, there are an infinite number of answers such as (0, 4), (2,2), (1/2, 3 1/2), (-6, 10) all of these sets will work in the first equation and the second

OpenStudy (anonymous):

thanks